OpenStudy (anonymous):

solve the equation 28=-4(n+15) n=???

mathslover (mathslover):

\[\huge{28=-4(n+15)}\] \[\huge{28=-4(n)-4(15)}\] Note : I applied The distributive property under addition that states that : \[\huge{a(b+c)=ab+ac}\] \[\huge{28=-4n-50}\] \[\huge{28+50=-4n}\] \[\huge{78=-4n}\]

mathslover (mathslover):

Can u solve further ?

OpenStudy (campbell_st):

divide both sides by -4 -7 = n + 15 i'll leave you to finish the question

mathslover (mathslover):

u can also do like @campbell_st did

OpenStudy (anonymous):

ok thanks working on it now

OpenStudy (anonymous):

ok so I divided both side and I got 3.75 and 1.75 then what divide?

OpenStudy (anonymous):

or is that not right?

OpenStudy (anonymous):

-19.5? @mathslover

jimthompson5910 (jim_thompson5910):

When you divide both sides by -4, you get -7 = n + 15 So now you need to solve -7 = n + 15 What do you get when you solve that last equation for n?

OpenStudy (anonymous):

-22 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Yayyy ! Thanks

OpenStudy (anonymous):

everybody

jimthompson5910 (jim_thompson5910):

It looks like once you have something like -7 = n + 15, you can see/get the solution very easily So I would practice with multiplying/dividing both sides like what is done above to sharpen your skills here.

jimthompson5910 (jim_thompson5910):

Ie practice going from 28=-4(n+15) to -7 = n + 15

OpenStudy (anonymous):

ok thanks for the advice! Im going to do another question here in a sec Id like if you could help

jimthompson5910 (jim_thompson5910):

alright, go ahead and post it